mario's recipe uses 2 cups of flour and makes 12 muffins. how many flour does he need to make 32 muffins?
Answer by Shin123(626)

Since 2 cups of flour makes 12 muffins, 2/3 cup of flour makes 4 muffins. Multiply 4*8=32 and 2/3*8=16/3=5 1/3 cups of flour.